OpenStudy (anonymous):

Which of the following is one of the factors of x2 + 4x – 32?

OpenStudy (anonymous):

you have to find two numbers whose product is -32 and whose sum is 4 .

OpenStudy (anonymous):

(x-a)(x-b) where a and b are the numbers you've find in step 1

OpenStudy (anikate):

jst divide everything by 2. Keep doing it till the end

OpenStudy (anonymous):

@Anikate ... ???

OpenStudy (anonymous):

x=-8 x=4

OpenStudy (anonymous):

so the answer is x-8? because my choices are x-8, x+4, x-4, and x-16

OpenStudy (anonymous):

??

OpenStudy (anonymous):

8+(-4) = 4, 8*(-4) =-32 so .... (x+8)(x-4)

OpenStudy (anonymous):

when you equate this to 0 x2 + 4x – 32? x=-8 x=4 which is the same as (x+8)(x-4),when you move it to the left

OpenStudy (anonymous):

the question was " ... one of the factors of x2 + 4x – 32?"

OpenStudy (anonymous):

good :)